OpenStudy (anonymous):

If y varies directly as x, and y = 16 as x = 4, find y for the x-value of 2. Answer 5 6 7 8

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

first do y=kx solve for k, the variation and then use that to find y when x=2 16=4k

OpenStudy (anonymous):

Got 6/

OpenStudy (anonymous):

just 6

OpenStudy (anonymous):

What is k?

OpenStudy (anonymous):

x is k just a letter replacement

OpenStudy (anonymous):

k=16/4 k=?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Yes, now use that to find y when x=2 y=4*2 y=?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

so 8 would be the answer? thanks!

OpenStudy (anonymous):

Correct. You're welcome